### v1

@@ 1 Introduction
In document processing research, Optical Character Recognition (OCR) has proven essential for transforming scanned documents and images into machine-readable text, facilitating further analysis. Initially, statistical methods [1] were used alongside OCR to extract information, followed by machine learning approaches. Subsequently, deep learning techniques [2], especially methods related to natural language processing (NLP), became crucial in advancing document understanding. Today, the focus has shifted towards Large Language Models (LLMs) [3], which, with their exceptional ability to model natural language in complex contexts, have further enhanced document comprehension and the automation of information extraction from extensive volumes of text.
@@
  • Document Image Classification: classifies document images into types such as invoices, scientific papers, and receipts [4].
  • Layout Analysis: examines a document’s structure, identifying elements like text, images, and tables [5];
  • Visual Information Extraction: extracts entities and relationships from unstructured content, considering text, visual elements, and layout [6];
  • Visual Question Answering: answers natural language questions based on a document’s content [7];
The two main motivations for building the BoundingDocs dataset [^1] The dataset is publicly available at https://huggingface.co/datasets/letxbe/BoundingDocs., that is focused on Information Extraction and Question Ansering, are:
@@
Current datasets do not effectively incorporate positional data, which is essential for reducing hallucinations and improving performance by enabling LLMs to understand document layout more precisely.
@@
  • RQ3: Does including layout information in prompts (e.g. [8, 9]) improve the model’s performance on document comprehension tasks?
@@ 2 State of the art
We provide an overview of the main models and techniques proposed for QA and Visual Question-Answering (VQA) [7]. We also discuss the features of the main datasets in the Document AI that we considered in our research.
@@ 2.1 Related Datasets
As summarized in Table 1, we selected datasets that best match our focus on comprehensive document understanding and advanced VQA, addressing challenges across both single-page and multi-page documents. For a more detailed review of datasets specific to Document Layout Analysis, please refer to our additional survey [10], which includes more datasets focused on layout-related tasks.
@@
VRDU [11] 2,556 Yes 1 0 1 1
Deepform [12] 60,000 Yes 1 1 1 1
DUDE [13] 4,974 Yes 1 1, 2, 4 1 1
FATURA [14] 10,000 Yes 2 5 2 1
SP-DocVQA [7] 12,767 No 1 3 1 1
MP-DocVQA [15] 5,929 No 1 2 1 1
FUNSD [16] 199 Yes 1 0 1 1
Kleister Charity [17] 2,788 No 3 1, 2 1 1
Kleister NDA [17] 540 No 3 1, 2 1 1
SROIE [18] 1,000 No 1 0 1 1
XFUND [19] 1,393 Yes 1 0 1 2,3,4,5,6,7,8
SynthTabNet [20] 600,000 Yes 1 5 2 1
CORD [21] 1,000 Yes 2 0 1 9
GHEGA [22] 246 Yes 1 0 1 10
@@
Among the foundational datasets, DocVQA [7, 15] stands as one of the earliest benchmarks dedicated to VQA on document images, focusing on understanding both textual and layout aspects of documents. Launched in 2020, DocVQA comprises multiple tasks designed to push the boundaries of document comprehension. The primary tasks include answering questions about individual document pages and analyzing multi-page documents — a crucial capability for real-world applications. The Single Page [7] subset of this dataset includes 50,000 questions over 12,767 documents, while the Multi Page [15] subset contains 46,436 questions spanning 5,929 documents (covering 47,952 pages in total). These datasets require models to interpret the visual structure of documents and to derive insights that go beyond simple text extraction.
@@
DUDE [13] builds on this foundational work by extending VQA to multi-domain, multi-purpose documents. The dataset provides 5,000 annotated PDF files with 18,700 question-answer pairs across various domains and time frames, making it a unique resource for tasks that merge Document Layout Analysis with complex, layout-based question answering. Unlike typical QA datasets, DUDE often requires multi-step reasoning, handling both content and structural queries. For instance, questions may include layout-based prompts such as “How many text columns are there?” or require arithmetic and comparison skills, presenting a challenging dataset for models trained primarily on text-based QA.
Another significant resource is Docmatix [23], developed by HuggingFace and released during our research period. Docmatix introduces a vast dataset with $2.4$ million images and $9.5$ million question-answer pairs from $1.3$ million PDF documents, making it one of the largest publicly available DocAI resources. Generated from the PDFA dataset, Docmatix uses OCR-extracted text to produce diverse QA pairs via an automated approach, offering comprehensive coverage of document types and layouts. This dataset provides only document images with paired QA responses, excluding the original OCR text, which shifts focus toward layout and image-based comprehension.
@@
In addition to these datasets, several others serve as standard benchmarks and are worth mentioning briefly. VRDU [11] includes two corpora—registration forms from the U.S. Department of Justice and ad-buy forms from the FCC—representing templates of varying complexity. The FATURA [14] dataset provides 10,000 images across 50 templates with imbalanced distributions for fields commonly found in invoices, such as buyer information and total amount, along with bounding box annotations for structured data extraction. Kleister [17] datasets offer specialized financial reports and legal documents, with Kleister Charity and Kleister NDA addressing entity extraction for key attributes. Deepform [12] offers approximately 20,000 labeled receipts for political ad purchases with labeled fields for specific political advertising details.
@@
Finally, FUNSD [16] and XFUND [19] are form-centric datasets focused on entity linking and key-value extraction in noisy, often multilingual documents. FUNSD includes 199 annotated forms in English, designed for form understanding, while XFUND broadens this to a multilingual setting with documents in seven languages, capturing the diversity of form structures globally.
@@ 2.2 Related methods
In recent years, the QA task [7, 15] has been approached in many ways, leveraging different techniques and model architectures. These methods can be broadly categorized into NLP-based, LLMs, and multimodal architectures, each addressing different aspects of document understanding and question answering.
NLP-based approaches build on general Question-Answering models, primarily focusing on text semantics without explicitly incorporating document layout or visual features. A prime example is BertQA [7], which utilizes a BERT architecture followed by a classification head to predict the start and end indices of an answer span. Modifications such as changes in hyperparameters and the introduction of new pre-training tasks have been explored in multiple works [24, 25], resulting in improved outcomes.
LLM-based methods leverage large language models to perform document understanding tasks by encoding structural and layout information directly into the input. For instance, LMDX [26] incorporates layout information via bounding box coordinates in the prompt, enhancing retrieval precision and reducing hallucination. DocLLM [27], which builds on the LayoutLM family, includes a specialized pretraining phase focused on structured layout data to improve document layout understanding. In contrast, NuExtract [28] is designed for extracting structured JSON data from documents, using training data derived from the Colossal Clean Crawled Corpus [29].
Multimodal architectures combine visual and textual features to enhance document comprehension across layout, content, and structure. Among OCR-free methods, mPLUG-DocOWL 1.5 [30] integrates a Vision Transformer (ViT) [31] with an LLM for comprehensive Document AI analysis, aligning layout and textual cues effectively without requiring separate OCR stages. Similarly, Donut [32] and Dessurt [33] operate without OCR preprocessing, directly integrating image and text data for robust document understanding.
In contrast, OCR-dependent models further refine document comprehension by incorporating OCR-based tokens. Hi-VT5 [15], for example, combines OCR tokens with visual features, optimizing its effectiveness for Question-Answering tasks that rely on precise textual information. Additionally, LayoutLMv3 [34] introduces visual patch embeddings in place of traditional CNNs to better align text, layout, and visual cues, resulting in improved performance on tasks requiring fine-grained structural interpretation.
@@ 3 Dataset construction
Figure 1: Dataset construction pipeline. The rectangles represent processes while the parallelograms represent foundation LLM models.
We base our new dataset, BoundingDocs, on the following datasets selected from Table 1: SP-DocVQA, MP-DocVQA, DUDE, Deepform, VRDU, FATURA, Kleister Charity, Kleister NDA, FUNSD, and XFUND. This collection encompasses a diverse range of document types, linguistic features, and question-answer formats, providing essential resources for training and evaluating advanced Document AI models.
@@ 3.1 Dataset format definition
For each document, a JSON file contains the annotation (examples in Figure 2). Each word in the answer is linked to its corresponding bounding box. Following established practices in the literature (e.g., LayoutLM [34], BERT [2]), the bounding boxes are normalized integers ranging from 0 to 1000 relative to the actual page size. Each bounding box is defined by a list of four values: the width, the height, the $X$ and $Y$ coordinates of the top-left vertex of the rectangle.
@@ 3.2 Producing annotations
A significant challenge comes from integrating various types of annotations into a unified structure. Datasets like Deepform, Kleister, and FATURA provide annotations that only establish a relationship between a key and its corresponding value in the text, such as annotating Address = 48 Woodford, SandyFord. However, these datasets lack essential positional information, such as the text’s location, frequency of occurrence, and page number. In contrast, datasets like VRDU and DocVQA include positional details that specify where the key value appears in the text. However, inconsistencies may arise because these datasets utilize different OCR tools, leading to variations in positional measurements and formats. To ensure consistent calculations for bounding box positions, Amazon Textract [35]. has been selected for this purpose.
@@ 3.2.1 Dataset preparation
Upon collecting and downloading the datasets, the following preliminary operations have been considered case by case. These additional steps are critical to standardize and prepare the datasets for the generation of annotations.
Annotation Conversion: When the annotations in a dataset have an undocumented or complex format, they are converted into a standardized, more straightforward format. This is particularly necessary for the VRDU dataset, where the original annotations require interpretation and conversion.
Filtering Pages/Questions: Some datasets contain redundant or irrelevant content, as unnecessary pages or questions, that have been removed. For instance, in the DocVQA dataset, pages from the Multi Page set were excluded from the Single Page set to prevent duplication. Additionally, for both DUDE and DocVQA datasets, pure visual questions, i.e., lacking the answer as recognized by the OCR in the image, are filtered out.
@@
Key-Value Association Creation: For certain datasets, key-value pairs for information extraction were manually generated from the annotations. For instance, in FUNSD and XFUND datasets the key-value associations are automatically created from existing document annotations. This step involves linking elements labeled as questions to their corresponding answers to facilitate coherent information extraction.
@@ 3.2.2 Matching annotations and OCR
To match the answer to each question with the data extracted by Textract [35], a script has been developed: the main challenge is to identify the correct word when the same value appears at multiple positions. A considerable time has been devoted to produce high quality annotations. This script, a significant part of our contribution, is used across all datasets with only slight modifications to match the different annotation formats.
For a document and a given key-value pair, where the key represents a label (such as “name,” “address,” or “date”) describing the type of information, and the value contains the actual data associated with that label, the script executes the following steps:
@@
    For each candidate line, verify that each word within it is also detected as a Word block by Textract and falls within the Line bounding box. These words and their positions form the extracted answer for each key.

### v2

@@ Abstract
footnotetext: Accepted for publication in the International Journal of Document Analysis and Recognition (IJDAR).
@@ 1 Introduction
In document processing research, Optical Character Recognition (OCR) has proven essential for transforming scanned documents and images into machine-readable text, facilitating further analysis. Initially, statistical methods [mihalcea-tarau-2004-textrank] were used alongside OCR to extract information, followed by machine learning approaches. Subsequently, deep learning techniques [DBLP:journals/corr/abs-1810-04805], especially methods related to Natural Language Processing (NLP), became crucial in advancing document understanding. Today, the focus has shifted towards Large Language Models (LLMs) [DBLP:journals/corr/abs-2303-08774], which, with their exceptional ability to model natural language in complex contexts, have further enhanced document comprehension and the automation of information extraction from extensive volumes of text.
@@
  • Document Image Classification: classifies document images into types such as invoices, scientific papers, and receipts [docimageclass];
  • Visual Information Extraction: extracts entities and relationships from unstructured content, considering text, visual elements, and layout [DBLP:conf/aaai/WangLJT0ZWWC21];
  • Visual Question Answering: answers natural language questions based on a document’s content [docvqa].
The two main motivations for building the BoundingDocs datasetThe dataset is publicly available at https://huggingface.co/datasets/letxbe/BoundingDocs., that is focused on Information Extraction and Question Answering, are:
@@
Current datasets do not effectively incorporate positional data, which is essential for reducing hallucinations and improving performance by enabling LLMs to understand document layout more precisely. In contrast, the proposed dataset, BoundingDocs, is specifically designed to capture positional information, which not only defines the dataset’s core feature but also serves multiple purposes. First, it allows for verifying whether a model correctly extracts both the value and its location, providing deeper insights into the model’s comprehension of document structure and helping to mitigate hallucinations. Second, it can be leveraged to enhance prompting by incorporating more detailed layout-aware instructions, further improving the model’s ability to interpret and generate structured outputs.
@@
  • RQ3: Does including layout information in prompts (e.g. [DBLP:journals/corr/abs-2306-00526, DBLP:conf/icdar/LamottWUSKO24]) improve the model’s performance on document comprehension tasks?
@@ 2 State of the art
We provide an overview of the main models and techniques proposed for Question Answering (QA) and Visual Question-Answering (VQA) [docvqa]. We also discuss the features of the main datasets in the Document AI that we considered in our research.
@@ 2.1 Related datasets
As summarized in Table 1, we selected datasets that best match our focus on comprehensive document understanding and advanced VQA, addressing challenges across both single-page and multi-page documents. One detailed review covering more datasets focused on layout-related tasks can be found in [gemelli2024].
@@
VRDU [vrdu] 2,556 Yes 1 0 1 1
Deepform [deepform] 60,000 Yes 1 1 1 1
DUDE [dude] 4,974 Yes 1 1, 2, 4 1 1
FATURA [fatura] 10,000 Yes 2 5 2 1
SP-DocVQA [docvqa] 12,767 No 1 3 1 1
MP-DocVQA [TITO2023109834] 5,929 No 1 2 1 1
FUNSD [funsd] 199 Yes 1 0 1 1
Kleister Charity [kleister] 2,788 No 3 1, 2 1 1
Kleister NDA [kleister] 540 No 3 1, 2 1 1
SROIE [sroie] 1,000 No 1 0 1 1
XFUND [xfund] 1,393 Yes 1 0 1 2,3,4,5,6,7,8
SynthTabNet [synthtabnet] 600,000 Yes 1 5 2 1
CORD [cord] 1,000 Yes 2 0 1 9
GHEGA [ghega] 246 Yes 1 0 1 10
@@
Among the foundational datasets, DocVQA [docvqa, TITO2023109834] stands as one of the earliest benchmarks dedicated to VQA on document images, focusing on understanding both textual and layout aspects of documents. Introduced in 2020, DocVQA comprises multiple tasks designed to push the boundaries of document comprehension. The main tasks include answering questions about individual document pages and analyzing multi-page documents — a crucial capability for real-world applications. The Single Page [docvqa] subset includes 50,000 questions over 12,767 documents, while the Multi Page [TITO2023109834] subset contains 46,436 questions spanning 5,929 documents (covering 47,952 pages in total). These datasets require models to interpret the visual structure of documents and to derive insights that go beyond simple text extraction.
@@
DUDE [dude] builds on this foundational work by extending VQA to multi-domain, multi-purpose documents. The dataset provides 5,000 annotated PDF files with 18,700 question-answer pairs across various domains and time frames, making it a unique resource for tasks that integrate Document Layout Analysis with complex, layout-based question answering. Unlike typical QA datasets, DUDE often requires multi-step reasoning, handling both content and structural queries. For instance, questions may include layout-based prompts such as “How many text columns are there?” or require arithmetic and comparison skills, presenting a challenging dataset for models trained primarily on text-based QA.
@@
In addition to these datasets, several others serve as standard benchmarks and are worth mentioning briefly. VRDU [vrdu] includes two corpora—registration forms from the U.S. Department of Justice and ad-buy forms from the FCC—representing templates of varying complexity. The FATURA [fatura] dataset provides 10,000 images across 50 templates with imbalanced distributions for fields commonly found in invoices, such as buyer information and total amount, along with bounding box annotations for structured data extraction. Kleister [kleister] datasets offer specialized financial reports and legal documents, with Kleister Charity and Kleister NDA addressing entity extraction for key attributes. Deepform [deepform] offers approximately 20,000 labeled receipts for political ad purchases with labeled fields for specific political advertising details.
@@
Finally, FUNSD [funsd] and XFUND [xfund] are form-centric datasets focused on entity linking and key-value extraction in noisy, often multilingual documents. FUNSD includes 199 annotated forms in English, designed for form understanding, while XFUND broadens this to a multilingual setting with documents in seven languages, capturing the diversity of form structures globally.
Recently, several unified datasets similar to BoundingDocs have been published, aiming to integrate multiple document sources or tasks sharing both similarities and differences with our approach. Docmatix [huggingface2024docmatix] and K2Q [DBLP:conf/emnlp/ZmigrodSSMNLV24] are Information Extraction datasets containing 2.1 million and 12,000 documents, respectively. Both generate question-answer pairs automatically using LLMs. Notably, K2Q, which is only available upon request, implements a similar strategy to ours by augmenting questions that were initially created using fixed templates at the dataset level. However, neither dataset includes information about the position of the answer within the text, which distinguishes BoundingDocs.
Additionally, recently published datasets such as BigDocs [DBLP:journals/corr/abs-2412-04626] and DocStruct4M [hu2024mplugdocowl] aim to aggregate multiple tasks beyond Information Extraction. These datasets contain 7 million multimodal documents and 4 million documents, respectively. While significantly larger than BoundingDocs, they fall outside our focus on business documents. Moreover, they encompass a wide range of tasks, including Screenshot2HTML, Table2LaTeX, ChartParsing, and TableParsing. Not all examples in these datasets follow a question-answering format, and, most notably, they do not provide bounding box information for the answers.
@@ 2.2 Related methods
In recent years, the QA task [docvqa, TITO2023109834] has been approached in many ways, leveraging different techniques and various model architectures. These methods can be broadly categorized into NLP-based, LLM-based, and multimodal architectures, each addressing different aspects of document understanding and question answering.
NLP-based approaches build on general Question-Answering models, primarily focusing on text semantics without explicitly incorporating document layout or visual features. A prime example is BertQA [docvqa], which utilizes a BERT architecture followed by a classification head to predict the start and end indices of an answer span. Modifications such as changes in hyperparameters and the introduction of new pre-training tasks have been explored in multiple works [Garncarek_2021, liu2019robertarobustlyoptimizedbert], resulting in improved outcomes.
LLM-based methods leverage large language models to perform document understanding tasks by encoding structural and layout information directly into the input. For instance, LMDX [perot2023lmdx] incorporates layout information via bounding box coordinates in the prompt, enhancing retrieval precision and reducing hallucinations. DocLLM [wang2023docllm], which builds on the LayoutLM family, includes a specialized pretraining phase focused on structured layout data to improve document layout understanding. In contrast, NuExtract [numind2024nuextract] is designed for extracting structured JSON data from documents, using training data derived from the Colossal Clean Crawled Corpus [dodge2021documentinglargewebtextcorpora].
Multimodal architectures combine visual and textual features to enhance document comprehension across layout, content, and structure. Among OCR-free methods, mPLUG-DocOWL 1.5 [hu2024mplugdocowl] integrates a Vision Transformer (ViT) [dosovitskiy2021imageworth16x16words] with an LLM for comprehensive Document AI analysis, aligning layout and textual cues effectively without requiring separate OCR stages. Similarly, Donut [kim2022ocrfreedocumentunderstandingtransformer] and Dessurt [davis2022endtoenddocumentrecognitionunderstanding] operate without OCR preprocessing, directly integrating image and text data for robust document understanding.
In contrast, OCR-dependent models further refine document comprehension by incorporating OCR-based tokens. Hi-VT5 [TITO2023109834], for example, combines OCR tokens with visual features, optimizing its effectiveness for Question-Answering tasks that rely on precise textual information. Additionally, LayoutLMv3 [huang2022layoutlmv3pretrainingdocumentai] introduces visual patch embeddings in place of traditional CNNs to better align text, layout, and visual cues, resulting in improved performance on tasks requiring fine-grained structural interpretation.
@@ 3 Dataset construction
Figure 1: Dataset construction pipeline. The process begins with two main dataset categories: QA datasets (red) and Key-Value extraction datasets (blue). Both categories are processed using AWS OCR (Textract), followed by annotation matching. For the Key-Value extraction datasets, questions are generated and rephrased using Mistral 7B v0.3. All processed components are then unified into BoundingDocs (purple).
We base our new dataset, BoundingDocs, on the following datasets selected from Table 1: SP-DocVQA, MP-DocVQA, DUDE, Deepform, VRDU, FATURA, Kleister Charity, Kleister NDA, FUNSD, and XFUND. Our selection focuses specifically on business document-based datasets, thereby excluding datasets such as InfographicsVQA [infographicvqa] and SlideVQA [DBLP:conf/aaai/TanakaNNHSS23], which primarily deal with non-business documents, as well as VisualMRC [DBLP:conf/aaai/TanakaNY21], which, despite being document-based, features abstractive rather than extractive answers, as detailed in Section 3.2.1.
Although most of the selected datasets originate from Key-Value extraction tasks, which may involve recurring question types, BoundingDocs ensures diversity by incorporating datasets with heterogeneous characteristics. Specifically, BoundingDocs includes documents with widely varying layouts, structural complexities, and languages, spanning multiple domains such as invoices, contracts, forms, receipts, and documents containing some handwritten-filled forms. This collection provides an essential resource for training and evaluating Document AI models.
@@ 3.1 Dataset format definition
For each document, a JSON file contains the annotation (examples in Figure 2). Each word in the answer is linked to its corresponding bounding box. Following established practices in the literature (e.g., LayoutLM [huang2022layoutlmv3pretrainingdocumentai], BERT [DBLP:journals/corr/abs-1810-04805]), the bounding boxes are normalized integers ranging from 0 to 1000 relative to the actual page size. Each bounding box is defined by a list of four values: the width, the height, the $X$ and $Y$ coordinates of the top-left vertex of the rectangle.
@@ 3.2 Producing annotations
A significant challenge comes from integrating various types of annotations into a unified structure. Datasets like Deepform, Kleister, and FATURA provide annotations that only establish a relationship between a key and its corresponding value in the text, such as annotating Address = 48 Woodford, SandyFord. However, these datasets lack essential positional information, such as the text’s location, frequency of occurrence, and page number. In contrast, datasets like VRDU and MP-DocVQA provide different types of positional information: the former provides bounding boxes for the values to be extracted, while the latter only specifies which page of the document contains the answer. However, inconsistencies may arise because these datasets utilize different OCR tools, leading to variations in positional measurements and formats. To ensure consistent calculations for bounding box positions, we selected Amazon Textract [amazon_textract], as it is both already deployed in our company’s document processing systems (reducing annotation costs and ensuring consistency) and one of the most widely used commercial OCR services.
@@ 3.2.1 Dataset preparation
Upon collecting and downloading the datasets the following preliminary operations have been considered case by case. These additional steps are critical to standardize and prepare the datasets for the generation of annotations. Note that when the test split is not public, as in Kleister, DUDE, and DocVQA, the documents in those test splits and their corresponding annotations are excluded from BoundingDocs.
Annotation Conversion: When the annotations in a dataset have a complex format, they are converted into a standardized, more straightforward format. This is particularly required for the VRDU dataset, where the original annotations require interpretation and conversion.
Filtering Pages/Questions: Some datasets contain redundant or irrelevant content, such as unnecessary pages or questions, which have been removed. For instance, in the DocVQA dataset, pages from the Multi Page set were excluded from the Single Page set to prevent duplication. Additionally, for both DUDE and DocVQA datasets, we filtered out all questions that can be defined as abstractive, meaning that the answer is not explicitly present in the document text but instead requires reasoning or synthesis. Since our objective is to provide the exact location of each answer within the document, it is essential to retain only extractive questions, where the answer can be directly found in the text. This ensures that every identified answer has a corresponding position in the document, which would not be possible for abstractive questions.
@@
Key-Value Association Creation: For FUNSD and XFUND, key-value pairs for information extraction were generated from the annotations. This step involves linking elements labeled as questions to their corresponding answers to facilitate coherent information extraction.
@@ 3.2.2 Matching annotations and OCR
To match the answer to each question with the data extracted by Textract [amazon_textract], a script has been developed whose main challenge is to identify the correct word when the same value appears at multiple positions. Our approach matches the annotated value with the extracted text and considers all occurrences as potential matches. While this method ensures broad coverage, it may lead to false positives when the same textual value appears in unrelated contexts. A thorough analysis of this aspect can be found in Section 3.3.
A considerable time has been devoted to produce high quality annotations. This script, a significant part of our contribution, is used across all datasets with only slight modifications to match the different annotation formats.
For a document and a given key-value pair, where the key represents a label (such as “name,” “address,” or “date”) describing the type of information, and the value contains the actual data associated with that label, the script executes the following steps (no brand new annotations are created):
@@
    For each candidate line that exceeds the Jaccard similarity threshold, we verify that each word of the ground truth answer is also detected as a Word block by Textract and falls within the Line bounding box. The set of all Word blocks (and their corresponding bounding boxes) that, when concatenated, reconstruct the original answer becomes the localized and annotated answer in BoundingDocs. If the answer is found in multiple locations on the page through this procedure, all occurrences are considered valid answers.
Note that all date annotations in the Kleister datasets follow the ’YYYY-MM-DD’ format. If a date appears in a different format within the document, its original annotation is standardized to this format. Our script is designed to handle this specific case through the use of regular expressions. This was not necessary for the other datasets.
3.2.3 Questions formulation
For datasets that are not originally designed for QA but only for key-value extraction, it is necessary, after matching the answers with the OCR output, to also generate the corresponding questions. Therefore, in addition to the three steps described in the previous section, the following two additional steps are performed: